Please help me.
What is the real value of x in the equation: log2(24) - log2(3) = log5(x)
thanks

Answer by Nate(3500)

log2(24) - log2(3) = log5(x)
log2(24/3) = log5(x)
log2(8) = log5(x)
log2(2^3) = log5(x)
3 = log5(x)
5^3 = x
125 = x